# A Multi-Stream Fusion Approach with One-Class Learning for Audio-Visual Deepfake Detection

### v2

@@ I Introduction
In this work, we propose a novel framework Multi-Stream Fusion Approach with One-Class Learning (MSOC) to tackle audio-visual deepfake detection, enhancing the generalization ability and interoperability. We extend the one-class learning approach, previously proposed in uni-modal contexts, to the audio-visual setting. We validate the generalization ability by resplitting the FakeAVCeleb [12] dataset and separating the unseen algorithms into the test set. We curated four test sets (RAFV, FAFV, FAFV, Unsynced) that cover all kinds of fake categories. We will make the dataset splits and model implementation publicly available upon the publication of this paper. Our contributions are summarized as:
@@ IV-A Dataset
Our Training set contains 350 videos each from categories RARV (Real), FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). Validation set contains 50 videos each from categories Real, FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). For Test set, we sampled 100 face swap (RAFV) and face swap-wav2lip (FVFA) videos not included in the training and validation sets. We generated 100 audio-only fake videos using a voice conversion library, named category FARV, due to FakeAVCeleb’s limited methods for audio fakes. It is important to note that our newly created FARV dataset is synchronized, whereas the FARV from the FakeAVCeleb[12] dataset is unsynchronized. Therefore, our FARV dataset has only one cue to detect a fake, making the unseen generation method more distinct. We also created a Unsynced category with 100 unsynchronized videos by shifting audio. Each of these four test datasets — RAFV, FVFA, FAFV, and Unsynced— consists of 100 real videos (RARV) and 100 unseen fake videos.
@@ IV-C Comparison Methods
We adopt some existing methods from the literature for comparison. The multimodal-dissonance model [18] utilizes a modality dissonance score (a distance between audio and visual features) to detect dissimilarities between modalities. AVDF [29] simply concatenates audio and visual features and maps them directly to audio-visual labels. The multilabel method [30] is trained using both audio and visual labels to address the issue that audio-visual labels may confuse the uni-modal feature extractors. MRDF_CE and MRDF_margin [9] utilize the cross- and within-modality regularization to maintain the unique features and differences of each modality during multimodal representation learning.
@@ IV-D Training Details
MRDF_CE [9] $54.38\pm 2.84$ $88.25\pm 0.83$ $47.25\pm 0.83$ $47.50\pm 0.61$
MRDF_Margin [9] $55.12\pm 1.02$ $86.88\pm 1.85$ $47.62\pm 1.19$ $47.88\pm 1.14$

### v3

@@ I Introduction
In this work, we propose a novel framework Multi-Stream Fusion Approach with One-Class Learning (MSOC) to tackle audio-visual deepfake detection, enhancing the generalization ability and interoperability. We extend the one-class learning approach, previously proposed in uni-modal contexts, to the audio-visual setting. We validate the generalization ability by resplitting the FakeAVCeleb [12] dataset and separating the unseen algorithms into the test set. We curated four test sets (RAFV, FAFV, FARV, Unsynced) that cover all kinds of fake categories. We will make the dataset splits and model implementation publicly available upon the publication of this paper. Our contributions are summarized as:
@@ IV-A Dataset
Our Training set contains 350 videos each from categories RARV (Real), FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). Validation set contains 50 videos each from categories Real, FARV, RAFV, and FAFV (excluding faceswap and faceswap-wav2lip). For Test set, we sampled 100 face swap (RAFV) and face swap-wav2lip (FAFV) videos not included in the training and validation sets. We generated 100 audio-only fake videos using a voice conversion library, named category FARV, due to FakeAVCeleb’s limited methods for audio fakes. It is important to note that our newly created FARV dataset is synchronized, whereas the FARV from the FakeAVCeleb[12] dataset is unsynchronized. Therefore, our FARV dataset has only one cue to detect a fake, making the unseen generation method more distinct. We also created a Unsynced category with 100 unsynchronized videos by shifting audio. Each of these four test datasets — RAFV, FAFV, FARV, and Unsynced— consists of 100 real videos (RARV) and 100 unseen fake videos.
@@ IV-C Comparison Methods
We adopt some existing methods from the literature for comparison. The multimodal-dissonance model [18] utilizes a modality dissonance score (a distance between audio and visual features) to detect dissimilarities between modalities. AVDF [29] simply concatenates audio and visual features and maps them directly to audio-visual labels. The multilabel method [30] is trained using both audio and visual labels to address the issue that audio-visual labels may confuse the uni-modal feature extractors. MRDF-CE and MRDF-Margin [9] utilize the cross- and within-modality regularization to maintain the unique features and differences of each modality during multimodal representation learning.
@@ IV-D Training Details
MRDF-CE [9] $54.38\pm 2.84$ $88.25\pm 0.83$ $47.25\pm 0.83$ $47.50\pm 0.61$
MRDF-Margin [9] $55.12\pm 1.02$ $86.88\pm 1.85$ $47.62\pm 1.19$ $47.88\pm 1.14$